Five customers deposit Rs. 5, 10, 15, 20 and 25 lakh in a fixed account with Islamic Bank respectively. The first three customers plan to deposit for seven years while the last two customers plan to deposit for five years. Since the fixed account is based on Mudaraba agreement, the Islamic Bank and they agreed on profit rate of 50:50 percent.
Suppose Islamic Bank earns a monthly profit of Rs. 50 lakhs by using this deposited money.

Required:
After giving a reasonable weightages to all customers, calculate the monthly profit rate of each customer.
1
Expert's answer
2020-06-15T11:13:22-0400

Monthly profit = Rs. 50 lakhs

Profit share = 50:50

Amount of profit distributed to customers = 0.5 *50 = Rs. 25 lakhs

Total deposit amount = Rs(5+10+15+20+25) = Rs.75 lakhs

Monthly profit =weighed proportion*Profit